# Two-Phase Channel Estimation for RIS-Assisted THz Systems with Beam Split

### v2

@@ III-C2 Common Coupled Angle Estimation at RIS
where $\mathbf{\Omega}_{k,\bar{m},l}$ and $\boldsymbol{\lambda}_{k,\bar{m},l}$ denote the sorted eigenbasis matrix and eigenvalue vector, respectively. Therefore, the matrix $\mathbf{\Omega}_{k,\bar{m},l}\triangleq[\mathbf{\Omega}_{k,\bar{m},l}^{\mathrm{S}}\,\,\mathbf{\Omega}_{k,\bar{m},l}^{\mathrm{N}}]$ can be divided into two parts: $\mathbf{\Omega}_{k,\bar{m},l}^{\mathrm{S}}\in\mathbb{C}^{N_{\mathrm{sub}}\times J_{k}}$ and $\mathbf{\Omega}_{k,\bar{m},l}^{\mathrm{N}}\in\mathbb{C}^{N_{\mathrm{sub}}\times\left(N_{\mathrm{sub}}-J_{k}\right)}$. The part $\mathbf{\Omega}_{k,\bar{m},l}^{\mathrm{S}}$ consists of the $J_{k}$ largest eigenvalues corresponding to the directions of largest variability, which spans the signal subspace. The remaining $N_{\mathrm{sub}}-J_{k}$ eigenvectors form $\mathbf{\Omega}_{k,\bar{m},l}^{\mathrm{N}}$, spans the noise subspace and is orthogonal to the signal subspace.
To calculate the MUSIC pseudo-spectrum, the practical implementation requires sampling the estimated angles. We define $\mathbf{\tilde{p}}_{k,\bar{m},l,q}\triangleq[\tilde{p}_{k,\bar{m},l,1},\dots,\tilde{p}_{k,\bar{m},l,Q_{\mathrm{B}}}]^{\mathsf{T}}$ as the MUSIC pseudo-spectrum, where $\tilde{p}_{k,\bar{m},l,q}$ is the sampling spectral value for $q=1,\dots,Q_{\mathrm{B}}$. To align these sampled values with our proposed algorithm, we adjust the angle sampling within the range of the CBS dictionary. Therefore, we define the sub-CBS dictionary $\mathbf{\Xi}_{\mathrm{R},\bar{m}}^{\mathrm{sub}}\in\mathbb{C}^{N_{\mathrm{sub}}\times Q_{\mathrm{B}}}$ as the first $N_{\mathrm{sub}}$ rows of the original CBS dictionary in (15).According to (17), each element in the sub-CBS dictionary $\xi_{\mathrm{R},\bar{m},q}^{\mathrm{sub}}$ can be derived as
$\xi_{\mathrm{R},\bar{m},q}^{\mathrm{sub}}\triangleq e^{-j2\pi\eta_{m}\left(\frac{q}{Q_{\mathrm{R}}}-1\right)\mathbf{n}_{\mathrm{sub}}},q=1,\dots,Q_{\mathrm{B}},$ (34)
@@
By mapping the sampled beams $\xi_{\mathrm{R},\bar{m},q}^{\mathrm{sub}}$ onto the noise space $\mathbf{\Omega}_{k,\bar{m},l}^{\mathrm{N}}$, the MUSIC algorithm constructs the pseudo-spectrum by evaluating the inverse of the projection’s norm at each potential direction, which can be given as
$\tilde{p}_{k,\bar{m},l,q}=\frac{1}{\left\|(\mathbf{\Omega}_{k,\bar{m},l}^{\mathrm{N}})^{\mathsf{H}}\xi_{\mathrm{R},\bar{m},q}^{\mathrm{sub}}\right\|_{2}^{2}},q=1,\dots,Q_{\mathrm{B}}.$ (35)
Given that any beam $\xi_{\mathrm{R},\bar{m},q}^{\mathrm{sub}}\in\mathbf{\Omega}_{k,\bar{m},l}^{\mathrm{S}}$ in the signal subspace should ideally have minimal projection onto the noise subspace $\mathbf{\Omega}_{k,\bar{m},l}^{\mathrm{N}}$, which means that the $\ell_{2}$-norm in (35) associated with the actual directions should approximate zero. The MUSIC algorithm typically relies on peak detection to find actual physical directions. To avoid undetectable marginal peaks, we extend the pseudo-spectrum by inserting the minimum value at its edges. The extended pseudo-spectrum can be express as
$\mathbf{p}_{k,\bar{m},l}=\left[\min\left(\tilde{\mathbf{p}}_{k,\bar{m},l}\right),\tilde{\mathbf{p}}_{k,\bar{m},l},\min\left(\tilde{\mathbf{p}}_{k,\bar{m},l}\right)\right].$ (36)
The potential directions can then be determined by employing local-maximum algorithms. However, due to the extended range of coupled angles, the estimated number of directions will be $K_{\text{cp}}$ times the original.
@@
$K_{\rm{cp},\max}=\lfloor\eta_{m}\left(2-1/Q_{\mathrm{R}}\right)\rfloor+1.$ (37)
@@
Recalling that the sampled beams $\xi_{\mathrm{R},m,q}^{\mathrm{sub}}$ can be written as
$\begin{split}\xi_{\mathrm{R},m,q}^{\mathrm{sub}}&=e^{-j2\pi\eta_{m}\left(\frac{q}{Q_{\mathrm{R}}}-1\right)\mathbf{n}_{\mathrm{sub}}}\\ &=e^{-j2\pi K_{\text{cp}}\mathbf{n}_{\mathrm{sub}}}e^{-j2\pi\eta_{m}\left(\frac{q}{Q_{\mathrm{R}}}-1\right)\mathbf{n}_{\mathrm{sub}}}\\ &=e^{-j2\pi\eta_{m}\left(\frac{q}{Q_{\mathrm{R}}}-1+\frac{K_{\text{cp}}}{\eta_{m}}\right)\mathbf{n}_{\mathrm{sub}}},\end{split}$ (38)
where $K_{\text{cp}}\in\mathbb{N},\forall q=1,\dots,Q_{\mathrm{B}}$. It is observed that when $q^{\prime}=q+K_{\text{cp}}Q_{\mathrm{R}}/\eta_{m}$, the element in sub-CBS dictionary will generate the identical beam, i.e., $\xi_{\mathrm{R},m,q}^{\mathrm{sub}}=\xi_{\mathrm{R},m,q^{\prime}}^{\mathrm{sub}}$. Therefore, the maximum number of identical beams $K_{\text{cp},\max}$ can be obtained by solving the integer equation $K_{\text{cp}}Q_{\mathrm{R}}/\eta_{m}=Q_{\mathrm{B}}$. By rearranging the terms and employing the relationship $Q_{\mathrm{B}}=2Q_{\mathrm{R}}-1$, we can obtain
$K_{\text{cp}}=\frac{\left(2Q_{\mathrm{R}}-1\right)\eta_{m}}{Q_{\mathrm{R}}}.$ (39)
Therefore, the maximum value of $K_{\text{cp}}$ is the integer part of this expression, which can be given as
$K_{\rm{cp},\max}=\left\lfloor\eta_{m}\left(2-\frac{1}{Q_{\mathrm{R}}}\right)\right\rfloor+1.$ (40)
According to Proposition 2, instead of the expected $J_{k}$ peaks in the MUSIC pseudo-spectrum, $K_{\text{cp},\max}J_{k}$ peaks will be detected. This makes the algorithm infeasible for distinguishing the true physical directions from fake ones within a single SC. Thankfully, it is observed from (37) that the spatial directions associated with true physical directions keep frequency-independent. Inspired by this, we propose a novel double sensing (DS) algorithm to acquire the true physical direction. In this approach, we select two SCs to calculate and compare their extended MUSIC pseudo-spectrum. The positions where peaks overlap indicate the true physical directions. However, due to the unavoidable noise and estimation errors, the pseudo-spectrum of different SCs may exist offsets. Therefore, in practical implement, we select the $J_{k}$ closest peaks to represent true physical directions. We introduce the peak distance $d$ to measure the distance between two peaks, which is given by
@@
To ensure the algorithm is both effective and efficient, it is crucial to select the appropriate SCs. Since $K_{\text{cp},\max}$ in (37) is affected by the choice of SCs, we should aim to satisfy $K_{\text{cp},\max}\leq 2$, which means each angle has at most two identical peaks. Employing the floor function inequality $\lfloor x\rfloor\leq x$ and substituting (5) into (37), we can obtain
$m\leq\frac{M}{2}+\frac{Mf_{c}Q_{\mathrm{R}}}{B\left(Q_{\mathrm{B}}\right)}+\frac{1}{2}.$ (42)
@@
$d\left(q_{m},q_{m^{\prime}}^{\prime}\right)=K_{\text{cp}}Q_{\mathrm{R}}f_{c}\left|\frac{1}{f_{m^{\prime}}}-\frac{1}{f_{m}}\right|.$ (43)
To distinguish true physical directions effectively, the peak distance should also be sufficiently large. According to (43), it is evident that the intervals between selected SCs should be large enough. Therefore, under the constraint in (42), we select the first and the $M/2$-th SCs to enhance the accuracy of estimation.
Fig. 5 depicts an example of estimated extended MUSIC pseudo-spectrum with the proposed DS approach. Assuming that the cascaded channel contains $J_{k}=3$ RIS-UE paths for an arbitrary UE with SNR $=20$ dB. It can be observed that the peaks overlap at actual physical directions, while those associated with fake spatial directions are completely separated. The estimated physical directions can be easily distinguished from the fake ones and accurately match the actual physical directions.
Figure 5: Normalized extended MUSIC pseudo-spectrum via DS-MUSIC algorithm.
III-D Channel Estimation for Other Subcarriers
Through the proposed two-step full CSI estimation, we obtained the cascaded channels for two subcarriers and the corresponding common physical angle. Based on this knowledge, the cascaded channel in (10) can be rewritten as
@@ III-D Channel Estimation for Other Subcarriers
where $\tilde{\mathbf{\Psi}}_{m}\!\triangleq\bar{\mathbf{R}}_{m}\left(\mathbf{A}_{\mathrm{T},m}^{*}\left(\hat{\boldsymbol{\varphi}}\right)\!\otimes\!\mathbf{A}_{\mathrm{R},m}\left(\hat{\boldsymbol{\gamma}}_{k}\right)\right)\!\in\!\mathbb{C}^{PT\times L^{2}J_{k}}$. The column size of the $\tilde{\mathbf{\Psi}}_{m}$ matrix is reduce from $Q_{\mathrm{B}}Q_{\mathrm{T}}$ of overcomplete dictionary to a multipath-related size $L^{2}J_{k}$. This significant reduction in the column size enhances both the efficiency and feasibility of the LS solution. Therefore, the angle-excluded coefficients $\tilde{\mathbf{x}}_{k,m}$ can be easily obtained as
@@
IV Analysis of Computational Complexity
In this section, we analyze the computational complexity of the proposed channel estimation protocol. We also compare our results with the other existing algorithms. Specifically, the complexity of the proposed scheme mainly stems from four major components:
@@ IV Analysis of Computational Complexity
    Cascaded channel estimation for two selected SCs by CBS-based GAMP approach: $\mathcal{O}\left(2PTQ_{\mathrm{T}}Q_{\mathrm{B}}\right)$.
@@
    Coupled directions acquisition at the RIS using DS-MUSIC algorithm: $\mathcal{O}\left(2N_{r}^{3}\right)$.
@@
$\mathcal{O}\left(2PTQ_{\mathrm{T}}Q_{\mathrm{B}}+2N_{r}^{3}+\left(M-2\right)L^{6}J_{k}^{3}\right).$ (48)
@@
For comparison, we also discuss the complexity of the conventional OMP scheme [30], BSA-OMP scheme [17] and CBS-GAMP scheme [23]. These algorithms perform under traditional channel estimation frameworks which means they need to be executed at each SC or simultaneously process all SCs. The computational complexity of conventional OMP algorithm is $\mathcal{O}\left(M{Q_{\mathrm{T}}}^{3}{Q_{\mathrm{R}}}^{6}\right)$. The BSA-OMP approach has the same complexity as the conventional OMP, $\mathcal{O}\left(M{Q_{\mathrm{T}}}^{3}{Q_{\mathrm{R}}}^{6}\right)$, since it similarly processes each subcarrier independently. The computational complexity of CBS-GAMP is $\mathcal{O}\left(MPTQ_{\mathrm{T}}Q_{\mathrm{B}}\right)$. Additionally, the number of SCs $M$ is typically large in wideband systems. Therefore, it is evident that the computational complexity of the proposed scheme is significantly lower than the other approaches mentioned.
V Numerical Results and Discussion

### v3

@@ IV-B Common Coupled Angle Estimation at RIS
where $\mathbf{\Omega}_{k,\bar{m},l}$ and $\boldsymbol{\lambda}_{k,\bar{m},l}$ denote the sorted eigenbasis matrix and eigenvalue vector, respectively. Therefore, the matrix $\mathbf{\Omega}_{k,\bar{m},l}\triangleq[\mathbf{\Omega}_{k,\bar{m},l}^{\mathrm{S}}\,\,\mathbf{\Omega}_{k,\bar{m},l}^{\mathrm{N}}]$ can be divided into two parts: $\mathbf{\Omega}_{k,\bar{m},l}^{\mathrm{S}}\in\mathbb{C}^{N_{\mathrm{sub}}\times J_{k}}$ and $\mathbf{\Omega}_{k,\bar{m},l}^{\mathrm{N}}\in\mathbb{C}^{N_{\mathrm{sub}}\times\left(N_{\mathrm{sub}}-J_{k}\right)}$. The part $\mathbf{\Omega}_{k,\bar{m},l}^{\mathrm{S}}$ consists of the $J_{k}$ largest eigenvalues corresponding to the directions of largest variability, which spans the signal subspace. The $N_{\mathrm{sub}}-J_{k}$ remaining eigenvectors, which form $\mathbf{\Omega}_{k,\bar{m},l}^{\mathrm{N}}$, span the noise subspace and are orthogonal to the signal subspace.
To calculate the MUSIC pseudo-spectrum, the practical implementation requires sampling the estimated angles. We define $\mathbf{\tilde{p}}_{k,\bar{m},l}\triangleq[\tilde{p}_{k,\bar{m},l,1},\dots,\tilde{p}_{k,\bar{m},l,Q_{\mathrm{B}}}]^{\mathsf{T}}$ as the MUSIC pseudo-spectrum, where $\tilde{p}_{k,\bar{m},l,q}$ is the sampling spectral value for $q=1,\dots,Q_{\mathrm{B}}$. To align these sampled values with our proposed algorithm, we adjust the angle sampling within the range of the CBS dictionary. Therefore, we define the sub-CBS dictionary $\mathbf{\Xi}_{\mathrm{R},\bar{m}}^{\mathrm{sub}}\in\mathbb{C}^{N_{\mathrm{sub}}\times Q_{\mathrm{B}}}$ as the first $N_{\mathrm{sub}}$ rows of the original CBS dictionary in (14). According to (15), each element in the sub-CBS dictionary $\boldsymbol{\xi}_{\mathrm{R},\bar{m},q}^{\mathrm{sub}}$ can be derived as
$\boldsymbol{\xi}_{\mathrm{R},\bar{m},q}^{\mathrm{sub}}\triangleq e^{-\jmath 2\pi\eta_{m}\left(\frac{q}{Q_{\mathrm{R}}}-1\right)\mathbf{n}_{\mathrm{sub}}},q=1,\dots,Q_{\mathrm{B}},$ (32)
@@
By mapping the sampled beams $\boldsymbol{\xi}_{\mathrm{R},\bar{m},q}^{\mathrm{sub}}$ onto the noise space $\mathbf{\Omega}_{k,\bar{m},l}^{\mathrm{N}}$, the MUSIC algorithm constructs the pseudo-spectrum by evaluating the inverse of the projection’s norm at each potential direction, which can be given as
$\tilde{p}_{k,\bar{m},l,q}=\frac{1}{\left\|(\mathbf{\Omega}_{k,\bar{m},l}^{\mathrm{N}})^{\mathsf{H}}\boldsymbol{\xi}_{\mathrm{R},\bar{m},q}^{\mathrm{sub}}\right\|_{2}^{2}},q=1,\dots,Q_{\mathrm{B}}.$ (33)
Therefore, the MUSIC pseudo-spectrum can be express as
$\mathbf{p}_{k,\bar{m},l}=\left[\tilde{p}_{k,\bar{m},l,1},\cdots,\tilde{p}_{k,\bar{m},l,Q_{\mathrm{B}}}\right].$ (34)
Given that any beam $\boldsymbol{\xi}_{\mathrm{R},\bar{m},q}^{\mathrm{sub}}\in\mathbf{\Omega}_{k,\bar{m},l}^{\mathrm{S}}$ in the signal subspace should ideally have minimal projection onto the noise subspace $\mathbf{\Omega}_{k,\bar{m},l}^{\mathrm{N}}$, which means that the $\ell_{2}$-norm in (33) associated with the actual directions should approximate zero. The MUSIC algorithm typically relies on peak detection to find actual physical directions. The potential directions can then be determined by employing local-maximum algorithms. However, due to the extended range of coupled angles, the estimated number of directions will be $K_{m,\text{cp}}$ times the original with $K_{\text{cp}}<K_{m,\text{cp,max}}$.
@@
$K_{m,\rm{cp},\max}=\lfloor\eta_{m}\left(2-1/Q_{\mathrm{R}}\right)\rfloor+1,$ (35)
[…]
@@
Recalling from (32) that the sampled beams $\boldsymbol{\xi}_{\mathrm{R},m,q}^{\mathrm{sub}}$ can be written as $\boldsymbol{\xi}_{\mathrm{R},m,q}^{\mathrm{sub}}=e^{-\jmath 2\pi\eta_{m}\left(\frac{q}{Q_{\mathrm{R}}}-1\right)\mathbf{n}_{\mathrm{sub}}}$. Based on the periodicity property of the complex exponential function, which satisfies $e^{-\jmath 2\pi K_{m,\text{cp}}}=1$ for any integer $K_{m,\text{cp}}\in\mathbb{N}$, $\boldsymbol{\xi}_{\mathrm{R},m,q}^{\mathrm{sub}}$ can be equivalently expressed as
$\begin{split}\boldsymbol{\xi}_{\mathrm{R},m,q}^{\mathrm{sub}}&=e^{-\jmath 2\pi K_{m,\text{cp}}\mathbf{n}_{\mathrm{sub}}}e^{-\jmath 2\pi\eta_{m}\left(\frac{q}{Q_{\mathrm{R}}}-1\right)\mathbf{n}_{\mathrm{sub}}}\\ &=e^{-\jmath 2\pi\eta_{m}\left(\frac{q}{Q_{\mathrm{R}}}-1+\frac{K_{m,\text{cp}}}{\eta_{m}}\right)\mathbf{n}_{\mathrm{sub}}}.\end{split}$ (36)
It is observed that when $q^{\prime}=q+K_{m,\text{cp}}Q_{\mathrm{R}}/\eta_{m}$, the element in sub-CBS dictionary will generate the identical beam, i.e., $\boldsymbol{\xi}_{\mathrm{R},m,q}^{\mathrm{sub}}=\boldsymbol{\xi}_{\mathrm{R},m,q^{\prime}}^{\mathrm{sub}}$. Therefore, the maximum number of identical beams $K_{m,\text{cp},\max}$ can be obtained by solving the integer equation $K_{m,\text{cp}}Q_{\mathrm{R}}/\eta_{m}=Q_{\mathrm{B}}$. By rearranging the terms and employing the relationship $Q_{\mathrm{B}}=2Q_{\mathrm{R}}-1$, we can obtain
$K_{m,\text{cp}}=\frac{Q_{\mathrm{B}}\eta_{m}}{Q_{\mathrm{R}}}.$ (37)
Therefore, the maximum value of $K_{m,\text{cp}}$ is the integer part of this expression, which can be given as
$K_{m,\rm{cp},\max}=\left\lfloor\eta_{m}\left(2-\frac{1}{Q_{\mathrm{R}}}\right)\right\rfloor+1.$ (38)
According to Proposition 2, instead of the expected $J_{k}$ peaks in the MUSIC pseudo-spectrum, $K_{\text{cp},\max}J_{k}$ peaks will be detected. This makes the algorithm infeasible for distinguishing the true physical directions from fake ones within a single SC. Thankfully, it is observed from (36) that the period $K_{m,\text{cp}}/\eta_{m}$ varies with SC frequency. Therefore, the periodic false spatial directions are frequency-dependent, while the true physical directions remain frequency-independent [20]. Inspired by this, we propose a novel double sensing (DS) algorithm to acquire the true physical direction. In this approach, we select two SCs to calculate and compare their MUSIC pseudo-spectrum, as these two SCs provide the necessary information to uniquely determine the coupled angles at RIS.
Assuming $\theta_{q_{\mathrm{p}}}$ to be the true physical direction, the corresponding sampled-beam $\boldsymbol{\xi}_{\mathrm{R},\bar{m},q_{\mathrm{p}}}$ can be determined according to (33). Then, the unique angular index $q_{\mathrm{p}}$ can be easily decoupled from (32), which is frequency-independent and remains consistent across all the SCs[20]. In contrast, the index of periodic false angles, given by $q^{\prime}_{\mathrm{p}}=q_{\mathrm{p}}+K_{m,\text{cp}}Q_{\mathrm{R}}/\eta_{m}$, varies among different SCs. Therefore, the positions where peaks overlap indicate the true physical directions. However, due to unavoidable noise and estimation errors, the pseudo-spectrum across different SCs may exhibit offsets. To estimate the true physical directions accurately, we first identify the $J_{k}$ closest peaks, which correspond to the overlapping peaks across different SCs. Therefore, we introduce the peak distance $d$ to quantify this deviation, given by:
@@
Selecting the closest peaks ensures consistency in the estimated directions and mitigates the impact of frequency-dependent variations. However, while the closest peaks are used for estimating the true directions, it is also essential to ensure that 1) the number of non-overlapping peaks are limited, and 2) these non-overlapping peaks are sufficiently separated. Satisfying these conditions allows effective discrimination between true and false physical directions. This can be achieved by selecting the appropriate SCs. Specifically, since $K_{m,\text{cp},\max}$ in (35) is affected by the choice of SCs, we should aim to satisfy $K_{m,\text{cp},\max}\leq 2$, which means each angle has at most two identical peaks. Employing the floor function inequality $\lfloor x\rfloor\leq x$ and substituting (5) into (38), we can obtain
$m<\frac{M}{2}+\frac{Mf_{c}}{BQ_{\mathrm{B}}}+\frac{1}{2}.$ (40)
@@
$d\left(q_{m},q_{m^{\prime}}^{\prime}\right)=K_{m,\text{cp}}Q_{\mathrm{R}}f_{c}\left|\frac{1}{f_{m^{\prime}}}-\frac{1}{f_{m}}\right|.$ (41)
To distinguish true physical directions effectively, the distance among non-overlapping peaks should also be sufficiently large. According to (41), it is evident that the intervals between selected SCs should be large enough. Therefore, under the constraint in (40), we select the first and the $M/2$-th SCs to enhance the accuracy of estimation [^7] Note that while using more subcarriers may offer additional spatial information, our two-subcarrier design leverages the deterministic beam split property for efficient peak alignment. Extending the alignment strategy to more than two subcarriers would require complex correspondence matching, which may introduce ambiguity and computational overhead without significant accuracy improvement..
Fig. 4 depicts an example of estimated MUSIC pseudo-spectrum utilizing the proposed DS approach. Assuming that the cascaded channel contains $J_{k}=3$ RIS-UE paths for an arbitrary UE with SNR $=20$ dB. It can be observed that the peaks overlap at actual physical directions, while those associated with fake spatial directions are completely separated. The estimated physical directions can be easily distinguished from the fake ones and accurately match the actual physical directions.
Figure 4: Normalized MUSIC pseudo-spectrum via DS-MUSIC algorithm.
V Channel Estimation for Other SCs
Through the proposed two-step full CSI estimation, we obtained the cascaded channels for two SCs and the corresponding common physical angle. Based on this knowledge, the cascaded channel in (10) can be rewritten as
@@ V Channel Estimation for Other SCs
where $\tilde{\mathbf{\Psi}}_{m}\!\triangleq\bar{\mathbf{R}}_{m}\left(\mathbf{A}_{\mathrm{T},m}^{*}\left(\hat{\boldsymbol{\varphi}}\right)\!\otimes\!\mathbf{A}_{\mathrm{R},m}\left(\hat{\boldsymbol{\gamma}}_{k}\right)\right)\!\in\!\mathbb{C}^{PT\times L^{2}J_{k}}$. The column size of the $\tilde{\mathbf{\Psi}}_{m}$ matrix is reduced from $Q_{\mathrm{B}}Q_{\mathrm{T}}$ of the overcomplete dictionary to a multipath-related size $L^{2}J_{k}$. Moreover, the incorporation of angular information transforms the observation matrix $\tilde{\mathbf{\Psi}}_{m}$ transformed into a full-rank matrix. Consequently, the sparse representation in (19) is reverted to the non-sparse frequency domain. This significant reduction in the column size and reconstruction of signal structure reformulate the channel estimation problem of remaining SCs a low-dimensional linear regression problem, enhancing both the efficiency and feasibility of the LS solution. Therefore, by employing LS algorithm, the angle-excluded coefficients $\tilde{\mathbf{x}}_{k,m}$ can be easily obtained as
@@
VI Analysis of Computational Complexity
In this section, we analyze the computational complexity of the proposed channel estimation protocol and compare it with that of existing algorithms. Specifically, the overall complexity of the proposed scheme is primarily attributed to the following four components:
@@ VI Analysis of Computational Complexity
    Cascaded channel estimation for two selected SCs by CBS-based GAMP approach: $\mathcal{O}\left(2PTQ_{\mathrm{T}}Q_{\mathrm{B}}N_{\mathrm{iter}}\right)$, where $N_{\mathrm{iter}}$ denotes the number of iterations required for GAMP to converge.
@@
    Coupled directions acquisition at the RIS using DS-MUSIC algorithm: $\mathcal{O}\left(2N_{\mathrm{R}}^{3}\right)$.
@@
$\mathcal{O}\left(2PTQ_{\mathrm{T}}Q_{\mathrm{B}}N_{\mathrm{iter}}+2N_{\mathrm{R}}^{3}+\left(M-2\right)L^{6}J_{k}^{3}\right).$ (46)
@@
For comparison, we also discuss the complexity of the conventional OMP scheme [35], BSA-OMP scheme [17] and CBS-GAMP scheme [26]. These algorithms perform under traditional channel estimation frameworks which means they need to be executed at each SC or simultaneously process all SCs. The computational complexity of conventional OMP algorithm is $\mathcal{O}\left(MPTQ_{\mathrm{T}}{Q_{\mathrm{R}}^{2}}\right)$. The BSA-OMP approach has the same complexity as the conventional OMP, $\mathcal{O}\left(MPTQ_{\mathrm{T}}{Q_{\mathrm{R}}^{2}}\right)$, since it similarly processes each SC independently. The computational complexity of CBS-GAMP is $\mathcal{O}\left(MPTQ_{\mathrm{T}}Q_{\mathrm{B}}N_{\mathrm{iter}}\right)$. Additionally, the number of SCs $M$ is typically large in wideband systems. Therefore, it is evident that the computational complexity of the proposed scheme is significantly lower than the other approaches mentioned.
VII Numerical Results and Discussion